OpenStudy (anonymous):

Write the slope-intercept equation of the line through the point (-2, 0) with slope 2.

OpenStudy (texaschic101):

y - y1 = m(x - x1) <-- point slope formula m (your slope) = 2 (-2,0) x1 = -2 and y1 = 0 now lets sub y - 0 = 2(x - (-2) y - 0 = 2(x + 2) y - 0 = 2x + 4 y = 2x + 4 any questions ?

OpenStudy (anonymous):

nope! thanks:)

OpenStudy (texaschic101):

no problem :)